The greatest acceleration or deceleration that a train may have a. The minimum time in  which the train may reach from one station to the other  separated by a distance  d is

Grade:11

2 Answers

Arun
4 years ago
 
The total distnace train has to cover is d.
The first half it will accelerate and the second half it’ll deaccelerate, and it’ll take half-half time for both acceleration and deacceleration.
 
Hence,
 
d/2 = a/2(t/2)2               (u = 0 ,because the train was at rest initially)
4d/a = t2
t = (4da)½.
Khimraj
4 years ago
Accn and declrcn each have taken time = t/2 S= 1/2(at/2)*t= at^2/4 t= 2√d/a .....................................